### v1

@@ 3.1 Simplest illustrative model
We first explain our proposed algorithm by applying it to the simplest workhorse economy with aggregate risk - the stochastic growth model of Brock and Mirman, (1972). This model can be accurately solved using conventional methods, so this section does not aim to claim a computational achievement of our method. Instead, it serves to lay out the proposed algorithm as transparently as possible, which is the easiest to do using a well-known work-horse model with well-known solution. In sections 4 and 5 we then apply our algorithm in settings that pose a substantial, if not prohibitive challenge for presently available numerical methods.
@@
where $\epsilon^{A}_{t}\sim\mathcal{N}(0,1)$. The firm rents capital $K_{t}$ and efficient units of labor $L$ on competitive spot markets and pays a rental rate $r^{K}_{t}=\alpha A_{t}K_{t}^{\alpha-1}L^{\alpha}$ on capital and a wage $w_{t}=(1-\alpha)A_{t}K_{t}^{\alpha}L^{-\alpha}$ for efficient units of labor. In each period, the representative household chooses how much to consume and how much to invest in capital. Capital evolves according to $K_{t+1}=(1-\delta)K_{t}+I_{t}$, where $I_{t}$ denotes the investment in capital. The households problem is given by
@@ 3.2.1 Main idea
At first sight, using the history of shock as an input may look like a step in the wrong direction for two reasons: first, for $T<\infty$ the sequence of shocks is only an approximately sufficient statistic for the state of the economy. Hence, by truncating the history, we introduce a limit to the precision that even a perfectly trained neural network could attain. Second, the dimensionality of $\mathbf{X}^{\text{seq},T}$ is sometimes larger than the dimensionality of $\mathbf{X}^{\text{state}}$. [^16] For example in this simple stochastic growth economy. Therefore, the sequence space approach would not work well together with approximation methods, for which a low-to-medium dimensional domain is pivotal, such as, for example, (adapative) sparse grids. [^17] See Krueger and Kubler, (2004) for sparse grids and Brumm and Scheidegger, (2017) for adaptive sparse grids. As we illustrate in this paper, however, the sequence approximation to the endogenous aggregate state is promising when used together with deep neural networks as function approximators.
@@ 3.2.2 Sufficiency of the truncated history of shocks
As an alternative to using lagged values of productivity $A_{t-T},A_{t-T+1},\dots,A_{t-1}$, we can use a truncated history of innovations to productivity $\epsilon_{t-T},\epsilon_{t-T+1},\dots,\epsilon_{t-1}$. This is because we have that
$\log(A_{t})=\rho\log(A_{t-1})+\sigma\epsilon_{t}=\rho(\rho\log(A_{t-2})+\sigma\epsilon_{t-1})+\sigma\epsilon_{t}$
$=\underbrace{g(\rho,\sigma,\epsilon_{t-T},\dots,\epsilon_{t})}_{\text{function of $\rho$, $\sigma$ and $T+1$ last innovations}}+\underbrace{\rho^{T}A_{t-T}}_{\text{truncation error}}.$ (9)
@@
The two eiqenvalues of the matrix $M$ are $\rho$ and $\alpha$ and hence the impact of $\log(A_{t-T})$ and $\log(K_{t-T})$ on $\log(K_{t})$ and $\log(A_{t})$ decays exponentially at rate $\max\{\rho,\alpha\}$. This also implies that in order to accurately predict the policy functions based on a truncated history of innovations in a setting where $\rho$ is close to 1, the history needs to be very long, leading to high dimensional state $\mathbf{X}^{\text{seq},T}_{t}$. [^18] With a slight abuse of notation, we use $\mathbf{X}^{\text{seq},T}$ to indicate the approximate state in sequence space based on either sequences of innovations ($\epsilon$) or values ($A$).
@@ 3.2.3 Remaining parts of the algorithm
The remaining parts of the algorithm follow Azinovic et al., (2022) and Azinovic and Zemlicka, (2024) and are summarized here for completeness.
@@
We train the neural network by minimizing a loss function. Specifically, we use the mean squared error in the household optimality condition, (5), expressed in terms of relative consumption error. By minimizing the (weighted) Euler equation error, we train the neural network policy function to take shape that is consistent with the optimal choices of the representative household.
@@
$ree(\mathbf{X}^{\text{seq},T}_{t},\mathbf{X}^{\text{state}}_{t},\bm{\rho})$ $:=\frac{(u^{\prime})^{-1}\left(\text{E}\left[\beta u^{\prime}(C_{t+1})(1-\delta+r_{t+1}^{K})\right]\right)}{C_{t}}-1$ (11)
@@
$\ell(\mathcal{D},\bm{\rho})$ $:=\frac{1}{N}\sum_{(\mathbf{X}^{\text{seq},T}_{t},\mathbf{X}^{\text{state}}_{t})\in\mathcal{D}}ree(\mathbf{X}^{\text{seq},T}_{t},\mathbf{X}^{\text{state}}_{t},\bm{\rho})^{2}.$ (18)
@@
$\bm{\rho}^{\text{new}}=\bm{\rho}^{\text{old}}+\alpha^{\text{learn}}\nabla_{\bm{\rho}}\ell(\mathcal{D},\bm{\rho})$ (19)
The ADAM optimizer modifies the equation (19) by introducing a parameter-specific adaptive learning rates rather than using a single constant learning rate for all parameters.
@@
To generate a dataset $\mathcal{D}$ of state-history pairs, we closely follow the ergodic simulation scheme used in the DEQN algorithm of Azinovic et al., (2022). We start the simulation with a random sample of initial exogenous and endogenous states, and the simulate exogenous states using their stochastic process, and evolve endogenous states using the current approximation of the policy function. Relative to the original DEQN method, our algorithm additionally keeps track of truncated shock histories over the simulation. This is important because we use the truncated histories of shocks as neural network inputs. A key advantage of using truncated shock histories as network input is that their distribution does not change over the training, as it is pinned-down by model primitives, i.e. shock distributions, rather than by policy functions.
@@ 3.3 Parameterization
We set relative risk aversion $\gamma=2$, depreciation $\delta=0.1$, and patience $\beta=0.95$. The capital share in production is set to $\alpha=\frac{1}{3}$, and the process for logarithm of TFP has persistence $\rho^{A}=0.8$ and $\sigma^{A}=0.03$. The model parameters are summarized in table 1.
Parameter $\gamma$ $\delta$ $\beta$ $\alpha$ $\rho^{A}$ $\sigma^{A}$
Meaning
relative
risk aversion
depreciation
of capital
patience
capital share
in production
persistence of
log TFP
std dev of
innovations log TFP
Value 2 0.1 0.95 $\frac{1}{3}$ 0.8 0.03
Table 1: Parameter values chosen for the Brock and Mirman, (1972) model.
@@ 3.4 Implementation
We train a densely connected feed-forward neural network to predict the savings rate in period t, $s_{t}$ as a function of the last $T$ realizations of innovations to aggregate productivity, i.e., $\mathbf{X}^{\text{seq},T}_{t}=[\epsilon_{t-T},\dots,\epsilon_{t-1},\epsilon_{t}]\in\mathbb{R}^{T}$ and $s_{t}=\mathcal{N}_{\bm{\rho}}(\mathbf{X}^{\text{seq},T}_{t})$. Approximating the savings rate as opposed to directly parameterizing the savings function $K_{t+1}$, has the advantage that we can ensure that savings rate is bounded in $(0,1)$ by applying a sigmoid activation in the output layer. [^20] The sigmoid function is defined as $\text{sigmoid}(x):=\frac{1}{1+e^{-x}}$. Because of that, our policy function is guaranteed to satisfy the budget constraint and also guaranteed to ensure non-negativity of consumption. Given the current capital stock and productivity level, we compute the quantity of resources on hand $M_{t}=A_{t}K_{t}^{\alpha}L^{1-\alpha}+(1-\delta)K_{t}$. The savings rate and resources on hand then imply the current consumption $C_{t}=(1-s_{t})M_{t}$ and the next periods capital $K_{t+1}=s_{t}M_{t}$. This structure is a good illustration of why our algorithm still needs to keep track of values of state variables. Even though states are not used as network input, they are still required to construct objects like resources at hand, marginal products, etc.
@@
obtain the new agg. state in seq. space: $\mathbf{X}^{\text{seq},T}_{j+1,i}=\left[[\mathbf{X}^{\text{seq},T}_{j,i}]_{2:T},\epsilon_{j+1,i}\right]$ (23)
obtain the new agg. state in state. space: $\mathbf{X}^{\text{state}}_{j+1,i}=[K_{j+1,i},A_{j+1,i}].$ (24)
Generating new training data in this way is computationally cheap and does not pose a bottleneck to our algorithm. This has the advantage that we can generate new training dataset after every episode, such that no datapoint is used twice during the training. Therefore, overfitting is not a concern. A key difference of our approach, relative to previous deep learning based approaches (e.g. Azinovic et al.,, 2022) is that the distribution of neural network inputs, $\mathbf{X}^{\text{seq},T}$, is exogenous and hence remains stationary throughout the training. The state-space representation $\mathbf{X}^{\text{state}}$ on the other hand, is converging to the ergodic set of states as the neural network learns the equilibrium policies.
@@ 3.5 Training
We use a history of $T=100$ previous productivity innovations as the approximate sufficient statistic replacing the state vector as a network input. We parameterize the history-based policy function using a densely-connected feed-forward neural network with three hidden layers. Each layer consists of 128 gelu activated neurons. [^21] The gelu activation function is given by $f(x)=x\Phi(x)$, where $\Phi(x)$ denotes the Gaussian cumulative distribution function. The output layer is sigmoid activated, ensuring that the predicted savings rate lies between 0 and 1. To compute the conditional expectation in the Euler equation we use Gauss-Hermite quadrature with $N^{\text{quad}}=8$ integration nodes. [^22] An alternative would be to discretize the AR(1) process, as in Rouwenhorst, (1995). For minimization of the loss function, we rely on the basic ADAM optimizer with a learning rate of $10^{-5}$ and mini-batches size of $256$. In each episode, our algorithm simulates $4096$ state-history pairs, implying $16$ gradient descent steps per episode. The hyperparameters are summarized in table 2.
Meaning
Quad.
nodes
Length of
hist. of shocks
# nodes
layer 1
# nodes
layer 2
# nodes
layer 3
# nodes
output layer
Value 8 100
128
128
128
1
Meaning
States per
episode
mini-batch
size
# episodes
Optimizer
learning
rate
Value 4096 256 40’000 Adam $10^{-5}$
Table 2: Hyperparameter values chosen for the neural network to solve the Brock and Mirman, (1972) model.
@@
Figure 1 shows the loss function during training. The loss function decreases steadily during the training and reaches a value well below $10^{-8}$ within a training run. [^23] The training run takes roughly 4 minutes on a Tesla T4 GPU.
Figure 1: Loss function when training the neural network for 50’000 episodes to solve the Brock and Mirman, (1972) model. The loss function is the mean squared error in the equilibrium conditions of the model, each episode consists of 4096 simulated states.
@@ 3.6 Accuracy
The error in the equilibrium conditions is an implicit measure, and depending on the model at hand, there might not be an obvious mapping between the magnitude of this implicit error measure and magnitude of actual policy error. For a complicated model, equilibrium conditions error might be the only available error measure. However, in this simple model, we have an access to high-quality reference solution provided by a standard grid-based solution methods. With a dense enough grid, the present model can be solved to very high accuracy, such that we can regard the obtained solution as a good proxy to the true policy function. Exploiting this classical solution, in figure 2, we show a comparison of the policy learned by the neural network, to the policy compute with standard policy time iteration (see, e.g., Judd,, 1998).
The comparison shows two things: first, the difference between the two policies is small, the mean relative difference is below $0.005\%$ and the 99.9th percentile is below 0.015%. This illustrates the the neural network is able to approximate the equilibrium policies to very high accuracy. Second, when comparing the explicit policy error (right panel) to the implicit error in the equilibrium conditions (left panel), we can observe that both errors are of the same magnitude. This indicates that the implicit error in the equilibrium conditions is a good proxy for the policy error. While this is a reassuring finding, it is important to point out that this finding depends on the model, so we can not generalize this conclusion to other models.
4 Application to OLG model
@@
In this section, we test the performance of our algorithm in more challenging setup. We compute the approximate equilibrium in an overlapping generations economy featuring portfolio choice, non-trivial market clearing, and more complex stochastic processes. In particular, we model 72 overlapping generations of households, who consume and save. Households allocate their savings into physical capital and risk-free bonds. While bonds are fully liquid, trading capital is subject to convex portfolio adjustment costs, and return to capital is exposed to aggregate risk. There are three sources of uncertainty: as in the Brock and Mirman, (1972) model, the logarithm of TFP follows an AR(1) process. Additionally, the economy switches between two regimes: normal times and disaster times. In disaster times, the depreciation to capital is stochastic and on average higher than during normal times. The regime process evolves according to a Markov transition matrix. Since capital is predominantly held by older households, the disaster leads to swings in the intergenerational wealth distribution.
@@
We train the neural network to predict equilibrium prices and policies as a function of truncated histories of aggregate shocks. For this economy, our algorithm keep track of the last $T$ realizations of aggregate shocks, which include $T$ innovations to the logarithm of TFP, $T$ innovations to the depreciation of capital, and the last $T$ regime realizations. The purpose of this section is to demonstrate that our sequence space method is capable of obtaining highly-accurate approximations to equilibria in economies featuring complex stochastic processes. The overlapping generations setup furthermore generates an obvious dependance of current outcomes on shock realizations, reaching far into the past.

### v2

@@ 3.1 Model
We first explain our proposed algorithm by applying it to the simplest workhorse economy with aggregate risk, the stochastic growth model of Brock and Mirman, (1972). This model can be accurately solved using conventional methods, so this section does not aim to claim a computational achievement of our method. Instead, it serves to lay out the proposed algorithm as transparently as possible. In sections 4, 5, and 6, we then apply our algorithm in settings that pose a substantial challenge for existing numerical methods.
@@
where $\epsilon^{A}_{t}\sim\mathcal{N}(0,1)$. The firm rents capital $K_{t}$ and efficient units of labor $L$ on competitive spot markets and pays a rental rate $r^{K}_{t}=\alpha A_{t}K_{t}^{\alpha-1}L^{\alpha}$ on capital and a wage $w_{t}=(1-\alpha)A_{t}K_{t}^{\alpha}L^{-\alpha}$ for efficient units of labor. In each period, the representative household chooses how much to consume and how much to invest in capital. Capital evolves according to $K_{t+1}=(1-\delta)K_{t}+I_{t}$, where $I_{t}$ denotes the investment in capital and $\delta$ denotes its depreciation rate. The households’ problem is given by
@@ 3.2.1 Main idea
At first sight, using the history of shocks as an input may look like a step in the wrong direction for two reasons: first, for $T<\infty$ the sequence of shocks is only an approximately sufficient statistic for the state of the economy. Hence, by truncating the history, we introduce a limit to the precision that even a perfectly trained neural network could attain. Second, the dimensionality of $\mathbf{X}^{\text{seq},T}$ is sometimes larger than the dimensionality of $\mathbf{X}^{\text{state}}$. [^19] For example in this simple stochastic growth economy. Therefore, the sequence space approach would not work well together with approximation methods, for which a low-to-medium dimensional domain is pivotal, such as, for example, (adaptive) sparse grids. [^20] See Krueger and Kubler, (2004) for sparse grids and Brumm and Scheidegger, (2017) for adaptive sparse grids. As we illustrate in this paper, however, the sequence approximation to the endogenous aggregate state is promising when used together with deep neural networks as function approximators.
@@ 3.2.2 Sufficiency of the truncated history of shocks
As an alternative to using lagged productivity levels, we could use a truncated history of innovations as the input to the neural network. [^21] Since the innovations are i.i.d. by construction, they may be easier for the network to process than highly persistent levels. Iterating the AR(1) implies
$\log(A_{t})$ $=\rho^{T}\log(A_{t-T})+\sigma\sum_{j=0}^{T-1}\rho^{j}\,\epsilon_{t-j}$ (9)
$=:g\!\left(\rho,\sigma,\epsilon_{t-T+1},\ldots,\epsilon_{t}\right)+\underbrace{\rho^{T}\log(A_{t-T})}_{\text{truncation error}}.$ (10)
@@
The two eigenvalues of $M$ are $\rho$ and $\alpha$, implying that the impact of $\log(A_{t-T})$ and $\log(K_{t-T})$ on $(\log(A_{t}),\log(K_{t}))$ decays geometrically at rate $\max\{|\rho|,|\alpha|\}$. This also implies that in order to accurately predict the policy functions based on a truncated history of innovations in a setting where $\rho$ is close to 1, the history needs to be very long, leading to high dimensional state $\mathbf{X}^{\text{seq},T}_{t}$. [^22] With a slight abuse of notation, we use $\mathbf{X}^{\text{seq},T}$ to indicate the approximate state in sequence space based on either sequences of innovations ($\epsilon$) or values (AA).
@@ 3.2.3 Remaining parts of the algorithm
The remaining parts of the algorithm follow Azinovic et al., (2022) (DEQN) and are summarized here for completeness.
@@
We train the neural network by minimizing a loss function. Specifically, we use the mean squared error in the household optimality condition, (5), expressed in terms of relative consumption error. By minimizing the (weighted) Euler equation errors, we train the neural network policy function to take shape that is consistent with the optimal choices of the representative household.
@@
$\text{ree}(\mathbf{X}^{\text{seq},T}_{t},\mathbf{X}^{\text{state}}_{t},\bm{\rho})$ $:=\frac{(u^{\prime})^{-1}\left(\text{E}\left[\beta u^{\prime}(C_{t+1})(1-\delta+r_{t+1}^{K})\right]\right)}{C_{t}}-1$ (12)
@@
$\mathcal{L}(\mathcal{D},\bm{\rho})$ $:=\frac{1}{N}\sum_{(\mathbf{X}^{\text{seq},T}_{t},\mathbf{X}^{\text{state}}_{t})\in\mathcal{D}}\text{ree}(\mathbf{X}^{\text{seq},T}_{t},\mathbf{X}^{\text{state}}_{t},\bm{\rho})^{2}.$ (19)
@@
$\bm{\rho}^{\text{new}}=\bm{\rho}^{\text{old}}+\alpha^{\text{learn}}\nabla_{\bm{\rho}}\mathcal{L}(\mathcal{D},\bm{\rho})$ (20)
The ADAM optimizer modifies the equation (20) by introducing parameter-specific adaptive learning rates rather than using a single constant learning rate for all parameters.
@@
To generate a dataset $\mathcal{D}$ of state-history pairs, we closely follow the ergodic simulation scheme used in the DEQN algorithm. We start the simulation with a random sample of initial exogenous and endogenous states, and then simulate exogenous states using their stochastic process, and evolve endogenous states using the current approximation of the policy functions. Relative to the original DEQN method, our algorithm additionally keeps track of truncated shock histories. This is important because we use the truncated histories of shocks as neural network inputs. A key advantage of using truncated shock histories as network input is that their distribution does not change during the training, as it is pinned-down by model primitives, i.e. shock distributions, rather than by policy functions.
@@ 3.3 Parameterization
We set relative risk aversion $\gamma=2$, depreciation $\delta=0.1$, and patience $\beta=0.95$. The capital share in production is set to $\alpha=\frac{1}{3}$, and the process for logarithm of TFP has persistence $\rho^{A}=0.8$ and $\sigma^{A}=0.03$. The model parameters are summarized in table A1.
@@ 3.4 Implementation
We train a densely connected feed-forward neural network to predict the savings rate $s_{t}$ in period t as a function of the last $T$ realizations of innovations to aggregate productivity, i.e., $\mathbf{X}^{\text{seq},T}_{t}=[\epsilon_{t-T},\dots,\epsilon_{t-1},\epsilon_{t}]\in\mathbb{R}^{T}$ and $s_{t}=\mathcal{N}_{\bm{\rho}}(\mathbf{X}^{\text{seq},T}_{t})$. Approximating the savings rate as opposed to directly parameterizing the savings function $K_{t+1}$, has the advantage that we can ensure that savings rate is bounded in $(0,1)$ by applying a sigmoid activation in the output layer. [^24] The sigmoid function is defined as $\text{sigmoid}(x):=\frac{1}{1+e^{-x}}$. Because of that, our policy function is guaranteed to satisfy the budget constraint and also guaranteed to ensure non-negativity of consumption. Given the current capital stock and productivity level, we compute the quantity of resources on hand $M_{t}=A_{t}K_{t}^{\alpha}L^{1-\alpha}+(1-\delta)K_{t}$. The savings rate and resources on hand then imply the current consumption $C_{t}=(1-s_{t})M_{t}$ and the next period’s capital $K_{t+1}=s_{t}M_{t}$. This illustrates why our algorithm still needs to keep track of values of state variables. Even though states are not used as network input, they are still required to construct objects like resources at hand, marginal products, etc.
@@
update the truncated history: $\mathbf{X}^{\text{seq},T}_{j+1,i}=\left[[\mathbf{X}^{\text{seq},T}_{j,i}]_{2:T},\epsilon_{j+1,i}\right]$ (24)
update the agg. state: $\mathbf{X}^{\text{state}}_{j+1,i}=[K_{j+1,i},A_{j+1,i}].$ (25)
Generating new training data in this way is computationally cheap and does not pose a bottleneck to our algorithm. This has the advantage that we can generate new training dataset after every episode, such that no datapoint is used twice during the training. Therefore, overfitting is not a concern. A key difference of our approach, relative to previous deep learning based approaches is that the distribution of neural network inputs, $\mathbf{X}^{\text{seq},T}$, is exogenous and hence remains stationary throughout the training. The state-space representation $\mathbf{X}^{\text{state}}$ on the other hand, is converging to the ergodic set of states as the neural network learns the equilibrium policies.
@@ 3.5 Training
We use a history of $T=100$ previous productivity innovations as the approximate sufficient statistic replacing the state vector as a network input. We parameterize the history-based policy function using a densely-connected feed-forward neural network with three hidden layers. Each layer consists of 128 gelu activated neurons. [^25] The gelu activation function is given by $f(x)=x\Phi(x)$, where $\Phi(x)$ denotes the Gaussian cumulative distribution function. The output layer is sigmoid activated, ensuring that the predicted savings rate lies between 0 and 1. To approximate the conditional expectation in the Euler equation we use Gauss-Hermite quadrature with $N^{\text{quad}}=8$ integration nodes. [^26] An alternative would be to discretize the AR(1) process, as in Rouwenhorst, (1995). For minimization of the loss function, we rely on the basic ADAM optimizer with a learning rate of $10^{-5}$ and mini-batch size of $256$. In each episode, our algorithm simulates $4096$ state-history pairs, implying $16$ gradient descent steps per episode. The hyperparameters are summarized in table A2.
@@
Figure 1 shows the loss function during training. The loss function decreases steadily during the training and reaches a value well below $10^{-8}$ by the end of training. [^27] The training run takes roughly 4 minutes on a Tesla T4 GPU.
Figure 1: Loss function when training the neural network for $\approx$40,000 episodes to solve the Brock and Mirman, (1972) model. The loss function is the mean squared error in the equilibrium conditions of the model, each episode consists of 4096 simulated states.
@@ 3.6 Accuracy
The error in the equilibrium conditions is an implicit measure, and depending on the model at hand, there might not be an obvious mapping between the magnitude of this implicit error measure and magnitude of actual policy error. For a complicated model, equilibrium conditions error might be the only available error measure. However, in this simple model, we have access to a high-quality reference solution provided by a standard grid-based solution methods. With a dense enough grid, the present model can be solved to very high accuracy, such that we can regard the obtained solution as a good proxy to the true policy function. Exploiting this classical solution, in figure 2, we show a comparison of the policy learned by the neural network, to the policy computed with standard policy time iteration (see, e.g., Judd,, 1998). The comparison shows two things: first, the difference between the two policies is small, the mean relative difference is below $0.005\%$, and the 99.9th percentile is below 0.015%. This illustrates that the neural network is able to approximate the equilibrium policies to very high accuracy. Second, when comparing the explicit policy error (right panel) to the implicit error in the equilibrium conditions (left panel), we can observe that both errors are of the same magnitude. This indicates that the implicit error in the equilibrium conditions is a good proxy for the policy error. Although this is a reassuring finding, it is important to point out that this finding depends on the model, so we cannot generalize this conclusion to other models.
4 Application to an OLG model
@@
In this section, we test the performance of our algorithm in a more challenging setup. We compute the approximate equilibrium in an overlapping generations economy featuring portfolio choice, non-trivial market clearing, and more complex stochastic processes for aggregate shocks. In particular, we model 72 overlapping generations of households who consume and save. Households allocate their savings into physical capital and risk-free bonds. While bonds are fully liquid, trading capital is subject to convex portfolio adjustment costs, and return to capital is exposed to aggregate risk. There are three sources of uncertainty: as in the Brock and Mirman, (1972) model, the logarithm of TFP follows an AR(1) process. In addition, the economy switches between two regimes: normal times and disaster times. In disaster times, the depreciation to capital is stochastic and on average higher than during normal times. The regime process evolves according to a Markov transition matrix. Since capital is predominantly held by older households, the disaster leads to swings in the intergenerational wealth distribution.
@@
We train the neural network to predict equilibrium prices and policies as a function of truncated histories of aggregate shocks. In this economy, our algorithm keeps track of the last $T$ realizations of each of the aggregate shocks, which include $T$ innovations to the logarithm of TFP, $T$ innovations to the depreciation of capital, and the last $T$ regime realizations. The purpose of this section is to demonstrate that our sequence space method is capable of obtaining highly-accurate approximations to equilibria in economies driven by multivariate stochastic processes. The overlapping generations setup furthermore generates an obvious long-range dependence of current outcomes on the realized sequence of past shocks.